Please solve -2x +6=-5
Answer by jessica43(140)

To solve this problem, you want to get x by itself on one side of the problem.
First, subtract 6 from each side of the equation:
-2x +6=-5
-2x + 6 - 6 = -5 - 6
-2x = -11
Now divide each side by -2 to get x by itself:
(-2x)/-2 = -11/-2
x = 5.5
So your answer is x = 5.5